# Low-Frequency Noise and Resistive Switching in $\beta$-Na0.33V2O5

### v1

@@ III RESULTS AND DISCUSSION
The rise in $V_{\text{th}}$ with decreasing temperature reflects the increasing stability of the CO polaronic insulating phase, which is disrupted under an applied electric field, leading to a transition into a more conducting phase and the emergence of strongly nonlinear current-voltage behavior.[20] This behavior is consistent with a field-induced delocalization of polarons and the collapse of the charge-ordered phase. The repeatable and non-destructive nature of the resistive transition, along with the absence of a required forming voltage during the initial application of high bias and the observed threshold-like behavior, suggests that the switching mechanism is not due to dielectric breakdown or filament formation.[59] Instead, it is more plausibly attributed to the field-induced destabilization of localized polarons into delocalized charge carriers, resulting in the breakdown of the CO state and the onset of band-like conduction.[60, 61] This transition is dominated by electronic effects, as opposed to thermal runaway due to Joule heating. This is supported by examining the power dissipated around the threshold voltage shown in Fig. 6 (b). Measurements of the thermal conductivity $\kappa$ of $\beta$-Na0.33V2O5 [26] indicate it increases as the sample cools below $T_{\text{CO}}$, improving the system’s ability to disperse heat into the environment through its coupling to the cryostat via the exchange gas within the sample space. Thus, more power would be required to locally heat and trigger the transition if Joule heating were responsible. Instead, we find that the power required to switch the device decreases with cooling.[62] Both theory and experiments suggest that the threshold switching behavior and the observed field values in the crystals are comparable to those found in other systems consisting of localized charge carriers undergoing hopping conduction, whether driven by polarons or disorder-induced Anderson localization.[60, 63] These systems include oxides such as Co1+xCr2-xO4, Co1-xLixO4, Pr1-xCaxMnO3, and numerous manganates, La2-xSrxNiO4, transition-metal-oxide glasses, and conjugated polymers, all of which exhibit threshold fields on the order of $10^{2}$$10^{6}$ V/cm.[57, 63, 58, 64, 62] Fig. 6 (c) shows that the resistance ratio $R_{\mathrm{OFF}}/R_{\mathrm{ON}}$ increases systematically with decreasing temperature, reflecting the progressive stabilization of the CO state in $\beta$-Na0.33V2O5. This enhancement in resistive contrast at low temperatures highlights the material’s potential for cryogenic memory applications, where a large and temperature-dependent $R_{\mathrm{OFF}}/R_{\mathrm{ON}}$ ratio is desirable for reliable state discrimination.

### v2

@@ III RESULTS AND DISCUSSION
The rise in $V_{\text{th}}$ with decreasing temperature reflects the increasing stability of the CO polaronic insulating phase, which is disrupted under an applied electric field, leading to a transition into a more conducting phase and the emergence of strongly nonlinear current-voltage behavior.[PARIJA20201166] This behavior is consistent with a field-induced delocalization of polarons and the collapse of the charge-ordered phase. The repeatable and non-destructive nature of the resistive transition, along with the absence of a required forming voltage during the initial application of high bias and the observed threshold-like behavior, suggests that the switching mechanism is not due to dielectric breakdown or filament formation.[Adda2022] Instead, it is more plausibly attributed to the field-induced destabilization of localized polarons into delocalized charge carriers, resulting in the breakdown of the CO state and the onset of band-like conduction.[Bryksin1998, Emin2006] This transition is dominated by electronic effects, as opposed to thermal runaway due to Joule heating. This is supported by examining the power dissipated around the threshold voltage shown in Fig. 6 (b). Measurements of the thermal conductivity $\kappa$ of $\beta$-Na0.33V2O5 [Markina2004] indicate it increases as the sample cools below $T_{\text{CO}}$, improving the system’s ability to disperse heat into the environment through its coupling to the cryostat via the exchange gas within the sample space. Thus, more power would be required to locally heat and trigger the transition if Joule heating were responsible. Instead, we find that the power required to switch the device decreases with cooling.[Lee2008] Both theory and experiments suggest that the threshold switching behavior and the observed field values in the crystals are comparable to those found in other systems consisting of localized charge carriers undergoing hopping conduction, whether driven by polarons or disorder-induced Anderson localization.[Bryksin1998, Ielmini2008] These systems include oxides such as Co1+xCr2-xO4, Co1-xLixO4, Pr1-xCaxMnO3, and numerous manganates, La2-xSrxNiO4, transition-metal-oxide glasses, and conjugated polymers, all of which exhibit threshold fields on the order of $10^{2}$$10^{6}$ V/cm.[Yamanouchi1999, Ielmini2008, Asamitsu1997, Basko2002, Lee2008] Fig. 6 (c) shows that the resistance ratio $R_{\mathrm{OFF}}/R_{\mathrm{ON}}$ increases systematically with decreasing temperature, reflecting the progressive stabilization of the CO state in $\beta$-Na0.33V2O5. This enhancement in resistive contrast at low temperatures highlights the material’s potential for cryogenic memory applications, where a large and temperature-dependent $R_{\mathrm{OFF}}/R_{\mathrm{ON}}$ ratio is desirable for reliable state discrimination.